# Abundance and SYZ conjecture in families of hyperkähler manifolds

### v1

@@ 3.2 $\mathbb{A}^{1}$-fibration on the semiample Teichmüller space
Proof: By 3.2 we may find a point $p^{\prime}\in F$ with parabolic $N^{1,1}_{p^{\prime}}$, and by the global Torelli theorem the preimage of $p^{\prime}$ in $\mathcal{T}^{\circ}(M,\ell)$ consists of the points indexed by $\pi_{0}(\mathcal{C}^{+}_{\ell,p^{\prime}})$. For any point $[I^{\prime}]$ in the preimage of $p^{\prime}$ the manifold $(M,I^{\prime})$ is not projective, so we have $[I^{\prime}]\in\mathcal{T}^{\circ}_{\mathrm{sa}}(M)$ by 3.1. Through each such point $[I^{\prime}]$ passes a unique degenerate twistor line that maps surjectively onto $F$. These lines are indexed by $\pi_{0}(\mathcal{C}^{+}_{\ell,p^{\prime}})$ and the latter is identified with $\pi_{0}(\mathcal{C}^{+}_{\ell,p})$ by the map $\chi_{p^{\prime}p}$ constructed above. The last claim follows from 3.2.
@@ 3.3 A global Torelli theorem for the semiample Teichmüller space
Proof: Part (i). If $[I_{1}]$ and $[I_{2}]$ are non-separated in $\mathcal{T}_{\mathrm{sa}}^{\circ}(M)$ then they are clearly non-separated in $\mathcal{T}^{\circ}(M)$. Conversely, let $[I_{1}]$ and $[I_{2}]$ be non-separated in $\mathcal{T}^{\circ}(M)$. Then they have the same period $p\in\mathcal{D}_{\ell}$. Let $U_{1}$ and $U_{2}$ be open neighbourhoods of $[I_{1}]$ and $[I_{2}]$ in $\mathcal{T}^{\circ}(M)$. Then the intersection $U_{1}\cap U_{2}$ is non-empty, and it contains a point $[I]\in\mathcal{T}^{\circ}(M,\ell)$ such that $(M,I)$ is non-projective. Such a point $[I]$ is contained in $\mathcal{T}_{\mathrm{sa}}^{\circ}(M)$ by 3.1. So $U_{1}\cap U_{2}\cap\mathcal{T}^{\circ}_{\mathrm{sa}}(M)$ is non-empty, hence $[I_{1}]$ and $[I_{2}]$ are non-separated in $\mathcal{T}_{\mathrm{sa}}^{\circ}(M)$.
@@ 3.4 The hyperkähler SYZ in families
Acknowledgements: We are grateful to Ekaterina Amerik for her interest and discussions.

### v2

@@ 3.2 $\mathbb{A}^{1}$-fibration on the semiample Teichmüller space
Proof: By 3.2 we may find a point $p^{\prime}\in F$ with parabolic $N^{1,1}_{p^{\prime}}$, and by the global Torelli theorem the preimage of $p^{\prime}$ in $\mathcal{T}^{\circ}(M,\ell)$ consists of the points indexed by $\pi_{0}(\mathcal{C}^{+}_{p^{\prime}})$. But since $N^{1,1}_{p^{\prime}}$ is parabolic, $\mathcal{C}^{+}_{p^{\prime}}=\mathcal{C}^{+}_{\ell,p^{\prime}}$ (see the proof of 3.1). For any point $[I^{\prime}]$ in the preimage of $p^{\prime}$ the manifold $(M,I^{\prime})$ is not projective, so we have $[I^{\prime}]\in\mathcal{T}^{\circ}_{\mathrm{sa}}(M)$ by 3.1. Through each such point $[I^{\prime}]$ passes a unique degenerate twistor line that maps surjectively onto $F$. These lines are indexed by $\pi_{0}(\mathcal{C}^{+}_{\ell,p^{\prime}})$ and the latter is identified with $\pi_{0}(\mathcal{C}^{+}_{\ell,p})$ by the map $\chi_{p^{\prime}p}$ constructed above. The last claim follows from 3.2.
@@ 3.3 A global Torelli theorem for the semiample Teichmüller space
Proof: Part (i). If $[I_{1}]$ and $[I_{2}]$ are non-separated in $\mathcal{T}_{\mathrm{sa}}^{\circ}(M)$ then they are clearly non-separated in $\mathcal{T}^{\circ}(M)$. Conversely, let $[I_{1}]$ and $[I_{2}]$ be non-separated in $\mathcal{T}^{\circ}(M)$. Then they have the same period $p\in\mathcal{D}_{\ell}$. Let $U_{1}$ and $U_{2}$ be open neighbourhoods of $[I_{1}]$ and $[I_{2}]$ in $\mathcal{T}^{\circ}(M)$ and $W=\mathcal{P}er(U_{1})\cap\mathcal{P}er(U_{2})\cap\mathcal{D}_{\ell}$. Then $W$ is an open neighbourhood of $p$ in $\mathcal{D}_{\ell}$, because the period map an open morphism. By 2.4 and 2.4, for a very general point $p^{\prime}\in W$ its preimage in $\mathcal{T}^{\circ}(M,\ell)$ consists of a single point $[I]$ such that $(M,I)$ is non-projective. By uniqueness of the preimage of $p^{\prime}$ we must have $[I]\in U_{1}\cap U_{2}$. The point $[I]$ is contained in $\mathcal{T}_{\mathrm{sa}}^{\circ}(M)$ by 3.1. So $U_{1}\cap U_{2}\cap\mathcal{T}^{\circ}_{\mathrm{sa}}(M)$ is non-empty, hence $[I_{1}]$ and $[I_{2}]$ are non-separated in $\mathcal{T}_{\mathrm{sa}}^{\circ}(M)$.
@@ 3.4 The hyperkähler SYZ in families
Acknowledgements: We are grateful to Ekaterina Amerik for her interest and discussions, and to Dominique Mattei for the useful remarks and comments.